OpenStudy (anonymous):

please help:) express 1/(√a)^4/3 in the form a^n, stating the value of n.

OpenStudy (anonymous):

\[1/(\sqrt{a})^{4/3}\]

OpenStudy (anonymous):

@hieuvo hey do you know how to work this one out too?:/

OpenStudy (anonymous):



OpenStudy (anonymous):

not the correct answer but thanks for your time:)

OpenStudy (anonymous):

it should be, otherwise, wrong question :)

OpenStudy (anonymous):

in the form a^n?:)

OpenStudy (anonymous):

ouch, sr, should be a^-1 :D

OpenStudy (anonymous):

theanswerin my text book is a^-23:/ weird

OpenStudy (anonymous):

so wierd :))

OpenStudy (anonymous):

im so confused:')

OpenStudy (anonymous):

\[\frac{ 1 }{ \sqrt{a}^{\frac{ 4 }{ 3 }} } = \frac{ 1 }{ a ^{\frac{ 2 }{ 3 }} } = a^{-2/3}\]

OpenStudy (anonymous):

@hieuvo it's 4/3 not 4/2